Find the Thevenin Equivaleut Circuit at the terminals shown. ki од Io R OB
Answer #1

* Vih: open circuit Voltage Vagi ki с А Is o :R Vih -V6-ki+ Ri = 0 Vm = i(rak) B ů= Io :: VT6 = Io (R-K) Volts. * Rai ThevninTheunin Equivalent Equivalent circuito A vod (R-k)... I (R- (R-K) CS Scanned with CamScanner